### v1

# GraphArena: Benchmarking Large Language Models on Graph Computational Problems

Abstract
The “arms race” of Large Language Models (LLMs) demands novel, challenging, and diverse benchmarks to faithfully examine their progresses. We introduce GraphArena, a benchmarking tool designed to evaluate LLMs on graph computational problems using million-scale real-world graphs from diverse scenarios such as knowledge graphs, social networks, and molecular structures. GraphArena offers a suite of 10 computational tasks, encompassing four polynomial-time (e.g., Shortest Distance) and six NP-complete challenges (e.g., Travelling Salesman Problem). It features a rigorous evaluation framework that classifies LLM outputs as correct, suboptimal (feasible but not optimal), or hallucinatory (properly formatted but infeasible). Evaluation of 10 leading LLMs, including GPT-4o and LLaMA3-70B-Instruct, reveals that even top-performing models struggle with larger, more complex graph problems and exhibit hallucination issues. Despite the application of strategies such as chain-of-thought prompting, these issues remain unresolved. GraphArena contributes a valuable supplement to the existing LLM benchmarks and is open-sourced at https://github.com/squareRoot3/GraphArena.
1 Introduction
As Large Language Models (LLMs) continue to evolve, accurately assessing their expanding capabilities becomes crucial but increasingly complex. Standardized tests, which are originally designed for humans, have become common LLM benchmarks in domains such as mathematics [30, 22, 50, 16], medicine [48, 15, 40], and multi-task collections [29, 65, 74, 89]. Despite their utility, concerns about data leakage in pretraining corpora [4, 82] cast doubt on whether LLMs are genuinely reasoning or merely memorizing. Crowd-sourced evaluations [88, 20], while reflective of real-world performance, are expensive and time-consuming [13]. Alternatively, evaluations using synthetic data like theorem proving [73, 80, 32], are rigorous but often lack real-world relevance.
To balance factors like cost, novelty, complexity, diversity, and practical applicability in LLM evaluation, the community is expanding its attention to modalities that go beyond text to include vision [50, 27], audio [84], tabular [66], and spatiotemporal [49] data. Among these, graphs stand out as a particularly promising avenue. Ubiquitous in many modern applications, graphs provide a robust framework for testing LLMs’ capabilities on interpret relational information, process non-sequential data, and generalize across diverse structures[44, 36, 59, 75, 18, 19]. Specifically, graph computational problems—which typically require systematic traversal or search algorithms to solve—serve as an excellent testbed for LLMs [25, 75]. Consider the task of finding the shortest path: a model must understand the entire graph structure, identify relevant nodes, logically deduce multiple steps to trace paths between nodes, and perform calculations to arrive at a correct conclusion.
Over the past year, several studies have assessed LLM performance on graph computational problems, including notable efforts like NLGraph [75], GraphQA [25], VisionGraph [45], GITA [79], and GraphInstruct [17]. While these studies have provided valuable insights, we identify three significant shortcomings in the existing benchmarks. First, they rely on synthetic graphs generated from models such as Erdős-Rényi [24] and Barabási-Albert [5], which may not accurately reflect real-world diversity. Second, these benchmarks focus predominantly on elementary graph tasks involving small-scale graphs, limiting tasks to simple queries and counts of basic elements like nodes, edges, triangles, and rings. More complex and challenging graph problems are largely overlooked. Third, they only require LLMs to provide straightforward answers—such as yes, no, or a numerical value—rather than detailed paths or reasoning processes. This format may be utilized by models through guesswork rather than through real comprehension of the underlying logic.
In this paper, we introduce GraphArena, a benchmarking tool designed to assess the reasoning capabilities of LLMs on addressing graph computational problems. Recognizing the limitations of existing benchmarks, GraphArena incorporates three core principles:
  • Realistic Graph Collection. GraphArena utilizes a rich assortment of million-scale real-world graphs, covering knowledge graphs, social networks, molecular structures, etc. We aim to maintain the integrity of the original graph topology and attributes throughout the sampling process.
  • Curated Task Selection. Our benchmark offers a diverse and challenging set of tasks that examine a broad spectrum of LLM reasoning skills. Illustrated in Figure 1, GraphArena includes 4 polynomial-time and 6 NP-complete challenges, which test both intuitive (System 1) and deliberative (System 2) cognitive processes [37]. These tasks require skills from simple computations to complex multi-step analytics, long-range location identification, and heuristic approximations.
  • Rigorous Evaluation Framework. GraphArena introduces a task-specific, fine-grained evaluation protocol to increase the rigor of assessments and reduce reliance on superficial pattern recognition. LLMs are required to identify the specific component or path in the graph that leads to the solution. Responses are then classified as correct, suboptimal (feasible but not optimal), hallucinatory (properly formatted but infeasible), or missing, enabling a nuanced comparison among LLMs.
Our comprehensive assessment of ten popular LLMs across 10,000 problems in GraphArena reveals substantial insights. Consistent with common perception, GPT-4o [2] and Llama3-70b-Instruct [3] stand out as the best closed-source and open-source models, respectively. However, even these leading models face challenges with NP-complete tasks and exhibit a propensity for hallucination, in which they frequently generate structurally correct but contextually infeasible responses. The hallucination issue is more severe on larger graphs and models with fewer parameters. While chain-of-thought prompting has shown some improvements, it remains insufficient to fully resolve the hallucination issue. Additionally, graph-aware LLMs fine-tuned on relevant corpora also fail to resolve this problem. These findings indicate that graph computational problems is a valuable testbed for evaluating and advancing LLMs towards the artificial general intelligence with strong reasoning capacities.
Figure 1: Overview of the GraphArena benchmark.
2 Benchmark Construction
Figure 1 provides an overview of our proposed GraphArena benchmark. In this section, we first discuss the collection and sampling methodologies for real-world graphs in Section 2.1, then offers detailed descriptions of each task within GraphArena and outlines the problem generation process in Section 2.2. Finally, Section 2.3 details the evaluation framework used to score the responses.
2.1 Dataset Collection
GraphArena distinguishes itself from previous benchmarks by utilizing real-world graphs instead of synthetic ones, offering a richer diversity of data. We collect graphs from the following five sources:
DBLP [39] is an academic database containing over 7 million publications under the CC0 1.0 Public Domain Dedication license. We use an undirected research collaboration graph for the period 2003-2018 from DBLP [72]. This graph consists of 1,354,852 nodes and 5,129,047 edges, with each node representing an author and each edge reflecting a collaboration that has occurred at least five times.
Social Network [63] is a repository comprising over 70 different social networks. We combine graphs from four large networks—Digg, Flixster, Lastfm, and Pokec—to form a dataset of 6,118,793 nodes and 40,647,227 edges. Each node symbolizes a user, and each edge denotes a friendship connection. All user names are pseudonymized for privacy. It is under the CC BY-SA 3.0 License.
DBpedia [11] extracts structured content from Wikipedia to form a comprehensive knowledge graph. We use DBpedia1M [81], a subset of the English version, which comprises 1,160,306 nodes and 7,214,631 edges. Each node represents an entity, and each edge denotes a relationship, with all original textual content retained as attributes. It is under the CC BY-SA 3.0 License.
Openflights [55] provides data for over 10,000 airports and their connecting flight routes. We construct an undirected graph where nodes are airports and edges are flight routes, weighted by the geographical distance between airports. We preserve the largest connected component in this graph, which consists of 3,390 nodes and 19,166 edges. It is under the DbCL v1.0 License.
PubChemQC [53] offers a quantum chemistry database detailing ground-state electronic structures. From the PCQM4Mv2 Dataset—part of the OGB Large-Scale Challenge [31] collected from PubChemQC—we incorporate 3,746,620 individual graphs. Each graph represents a molecule, where nodes are atoms and edges are chemical bonds, with atom types preserved as node attributes. It is under the CC BY 4.0 License.
We employ a random walk with restart strategy initiating from randomly selected nodes to effectively sample subgraphs. This technique enables us to identify the top-k most frequently visited nodes, forming local dense subgraphs that closely maintain the topological features of the original graphs. For the PubChemQC dataset, we sample molecular graphs directly by specified sizes. All graphs are treated as unweighted and undirected, except for OpenFlights which is weighted.
2.2 Task Design
Wikipedia lists 73 graph computational problems in graph theory, not counting variants. Selecting an appropriate subset for benchmarking is a notable challenge. Previous benchmarks often focused on simpler tasks like querying the existence and counts of basic graph elements such as nodes, edges, triangles, and rings [75, 25]. In contrast, we aim at incorporating more complex problems that demand advanced reasoning abilities.
To establish a nuanced evaluation of responses, we decompose each task into two requirements: satisfying the first requirement indicates a feasible solution, whereas fulfilling both requirements denotes a correct solution. We have chosen four polynomial-time tasks that require increasingly thorough understandings of graph structures:
  • Common Neighbor: For a graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$ with vertices $v_{1},v_{2}\in\mathcal{V}$, the task involves identifying the subset $S\subseteq\mathcal{V}$ where each $u\in S$ is connected to both $v_{1}$ and $v_{2}$. The objectives are to (1) identify these common neighbors, and (2) maximize their count. We utilize the DBLP academic collaboration graph for this task.
  • Shortest Distance: In a graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$, determine the shortest path from node $v_{1}$ to $v_{2}$. The goals are to (1) find a viable path $(v_{1},u_{1},\ldots,v_{2})$, and (2) minimize its hop count. We apply this task to the DBpedia knowledge graph.
  • Connected Component: In this task, the model identifies one representative node from each connected component within a graph. For a given graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$, the objectives are to (1) identify a set of vertices such that each is from a distinct connected component, and (2) ensure that these vertices represent all possible connected components of the graph. This task is applied to the Social Network dataset.
  • Graph Diameter: The diameter of a graph is the maximum distance between any pair of nodes in the graph. Given a graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$, this task entails finding (1) a shortest path between two arbitrary vertices and (2) ensuring it is the maximum possible among all shortest paths. This is performed using the DBpedia knowledge graph.
Each task is briefly described by a keyword (Neighbor, Distance, Component, Diameter) to facilitate easy reference. For the NP-complete challenges, we have selected six representative tasks:
  • Maximum Clique Problem (MCP): Identify the largest complete subgraph, or clique, in a given graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$. The task entails (1) finding a clique $\mathcal{C}\subseteq\mathcal{V}$ that is feasible within the graph, and (2) ensuring that $\mathcal{C}$ is the largest among all possible cliques. This is implemented using the DBLP dataset.
  • Maximum Independent Set (MIS): Select the largest set of mutually nonadjacent nodes from a graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$. The task is to (1) identify an independent set $\mathcal{S}\subseteq\mathcal{V}$, and (2) ensure that $\mathcal{S}$ is the largest among all feasible sets. This task is performed using the Social Network dataset.
  • Minimum Vertex Cover (MVC): Given a graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$, find a subset of nodes $\mathcal{S}\subseteq\mathcal{V}$ such that each edge in $\mathcal{G}$ has at least one endpoint in $\mathcal{S}$. The Minimum Vertex Cover Problem is to (1) find the vertex cover $\mathcal{S}$ and (2) ensure the size of $\mathcal{S}$ is minimum among all vertex covers of $\mathcal{G}$. We explore this using the Social Network dataset.
  • Maximum Common Subgraph (MCS): Compare two graphs $\mathcal{G}$ and $\mathcal{H}$ to find the largest common subgraph. The objectives are to (1) determine a node-induced common subgraph $\mathcal{S}$ between $\mathcal{G}$ and $\mathcal{H}$, and (2) maximize the size of $\mathcal{S}$. The PubChemQC dataset is used for this task.
  • Graph Edit Distance (GED): For two graphs $\mathcal{G}$ and $\mathcal{H}$, determine the minimum edit distance via node mappings. This involves (1) establishing a node mapping that aligns $\mathcal{G}$ with $\mathcal{H}$, and (2) minimizing the edit operations required. These operations include adding or deleting an edge or isolated node, or relabeling a node. The PubChemQC dataset is utilized here.
  • Traveling Salesman Problem (TSP): Solve the TSP in a weighted graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$, where $w:\mathcal{E}\rightarrow\mathbb{R}^{+}$ assigns a positive weight to each edge, representing the travel distance. The objectives are to (1) find a route $\mathcal{P}$ that visits each node exactly once and returns to the starting point, and (2) minimize the total travel distance. We utilize the OpenFlights dataset for this challenge. To ensure every pair of nodes is connected, thereby guaranteeing a feasible solution, we convert the dataset into a complete graph by adding edges to represent the shortest possible distances between nodes that are not directly connected.
Problem Generation: For each of the 10 tasks, we randomly sample 500 small and 500 large graphs to create two distinct subsets, yielding a total of 10,000 graphs. The size ranges for these graphs are tailored to the complexity of the tasks. For the Neighbor and Distance tasks, small graphs contain 4 to 19 nodes, while large graphs contain 20 to 50 nodes. For Component, Diameter, MCP, MIS, and MVC, small graphs have 4 to 14 nodes, and large graphs range from 15 to 30 nodes. For MCS, GED and TSP, small graphs have 4 to 9 nodes and large graphs range from 10 to 20 nodes.
After sampling graphs with the given size, we encode graph problem into text based on templates, a common approach in previous benchmarks [75, 17]. An example of this can be seen in the bottom left of Figure 1. The process begins with an introduction and definition of the task, followed by a randomly selected example with the correct answer to demonstrate the problem-solving process. Subsequently, the actual problem to be solved is listed, including detailed graph information and task requirements. While the graphs used may not appear large, their conversion into text often results in problems containing up to 6,000 tokens, posing a significant long-range challenge. Detailed statistics on graph sizes and problem lengths, along with examples for each task, are provided in the appendix.
2.3 Evaluation Framework
Unlike previous benchmarks that only verify responses as yes, no, or a numeric value, we develop a more detailed evaluation framework to prevent pattern-based guesses. For each task, LLMs are required to output the relevant graph component or the path that contributes to the final answer, as indicated by the green sections in each graph in Figure 1. Responses are categorized into four types through a checker: (1) Correct: The solution satisfies both specified task requirements; (2) Suboptimal: The solution satisfies the first requirement but not the second; (3) Hallucinatory: The solution is in the correct format but does not satisfy either requirement; (4) Missing: The LLM fails to provide a solution in the correct format.
For example, in the bottom middle of Figure 1, Deepseek-V2 correctly identifies the maximum clique of four authors. In contrast, GPT-4o finds a feasible but suboptimal solution with three nodes, while Llama3-8b produces a hallucinatory response, erroneously linking Claudiu and Fabien—a connection that does not exist. If we merely report numbers as in existing graph problem benchmarks, this error would be misclassified as a correct answer.
Metrics. Given the rarity of missing results in our experiments (less than 1% in most cases), our primary metrics focus on Accuracy (the proportion of correct answers), Feasibility (the proportion of both correct and sub-optimal answers), and Hallucination (the proportion of hallucinatory responses). We also introduce ranking-based metrics for additional comparison, particularly useful when most models yield suboptimal solutions. These metrics include Mean Reciprocal Rank (MRR), Top-1 Probability, and Top-3 Probability.
3 Experiments
Experimental Setup. In our experiments, we evaluated ten popular LLMs in GraphArena, encompassing both closed-source and open-source models. This included single models and mix-of-experts with diverse scales. Among the closed-source models, we tested the latest GPT-4o (gpt-4o-2024-05-13) [54], the well-known GPT-3.5 (gpt-3.5-turbo-0125) [56], and the cost-effective Claude3-haiku [6]. In the open-source category, our assessments included the Llama series (Llama3-70b-Instruct and Llama3-8b-Instruct) from Meta, the Qwen series (Qwen1.5-72b-Chat and Qwen1.5-8b-Chat) from AliCloud, and Gemma-7b from Google Cloud [1]. Additionally, we evaluated two mix-of-expert LLMs: Deepseek-V2 [10] with 230 billion parameters (21 billion active) and Mixtral-7x8b [35] featuring 47 billion parameters (13 billion active).
For closed-source LLMs and open-source LLMs larger than 10 billion parameters, we utilized services from cloud providers, including OpenAI, Claude, AliCloud [23], Deepseek [60], and AIMLAPI [7]. For smaller models with fewer than 10 billion parameters, inferences were conducted on our local machines equipped with four NVIDIA H800 PCIe 80GB GPUs. EGiven the high costs and time-intensive nature of evaluating LLMs, we limited our runs to a single instance per model across 10,000 problems in GraphArena. To encourage consistency in responses, we used a low temperature setting of 0.1 and imposed no constraints on the output length of any model. All responses were documented for public reference.
3.1 Main Results
Table 1: Average rankings and performance of 10 LLMs on 4 Polynomial-time and 6 NP-complete tasks across small and large graphs. All metrics excluding MRR are scaled to [0, 100]. Acc., Fea., and Hallu. represent Accuracy, Feasibility, and Hallucination probability, respectively. For all metrics except Hallucination, higher values indicate better performance. The best-performing open and closed-source models are highlighted in bold.
Polynomial-Time Tasks (Small) Polynomial-Time Tasks (Large)
LLM MRR Top-1 Top-3 Acc. Fea. Hallu. MRR Top-1 Top-3 Acc. Fea. Hallu.
GPT-4o 0.82 78.00 81.53 71.90 82.30 17.45 0.63 55.95 63.29 40.65 59.00 39.55
GPT-3.5 0.46 36.58 42.66 34.00 52.75 47.25 0.30 18.37 25.82 14.20 30.30 69.55
Claude-haiku 0.53 45.51 49.08 43.10 53.90 46.00 0.34 23.89 29.63 18.85 33.40 66.50
Llama3-70b 0.72 66.42 70.80 61.20 72.70 27.30 0.54 45.58 53.93 31.55 50.95 49.05
Llama3-8b 0.40 30.74 34.98 28.55 45.30 53.90 0.23 12.45 15.85 9.40 19.65 78.20
Qwen1.5-72b 0.50 41.62 46.87 39.05 58.05 41.80 0.34 24.51 28.14 20.45 28.85 70.80
Qwen1.5-7b 0.23 12.79 14.67 11.85 20.75 79.10 0.16 6.11 7.27 4.60 8.40 91.20
Deepseek-V2 0.63 56.11 62.53 51.35 64.65 34.95 0.44 34.50 41.59 24.70 38.00 61.15
Mixtral-7x8b 0.48 39.81 45.04 37.80 53.00 46.35 0.31 19.24 28.00 13.85 31.75 67.20
Gemma-7b 0.35 26.29 29.39 25.15 38.60 61.40 0.20 10.31 12.21 9.20 16.05 83.95
NP-Complete Tasks (Small) NP-Complete Tasks (Large)
GPT-4o 0.62 52.36 62.84 40.83 74.20 23.23 0.42 28.68 43.28 5.93 49.17 45.60
GPT-3.5 0.47 33.28 46.91 25.17 73.20 26.33 0.26 11.32 22.81 2.27 38.57 59.77
Claude-haiku 0.56 43.50 58.09 33.47 78.00 21.10 0.33 17.98 34.67 3.93 40.90 58.13
Llama3-70b 0.58 47.50 60.47 36.77 75.47 24.53 0.37 21.01 40.22 4.80 48.77 50.07
Llama3-8b 0.37 25.15 33.40 20.17 58.07 41.73 0.22 7.11 21.39 1.37 31.17 68.77
Qwen1.5-72b 0.48 36.25 48.14 28.30 69.73 27.40 0.30 13.62 31.67 3.93 47.97 51.63
Qwen1.5-7b 0.28 16.00 22.67 12.70 50.17 49.63 0.18 4.85 12.43 1.13 25.63 73.83
Deepseek-V2 0.56 43.48 58.68 33.70 74.17 24.47 0.35 19.80 36.68 5.43 45.00 54.97
Mixtral-7x8b 0.33 19.76 28.15 15.57 59.43 37.43 0.21 7.99 15.52 2.13 24.20 75.50
Gemma-7b 0.31 16.95 26.69 12.93 57.13 42.60 0.21 7.17 16.20 1.67 29.40 64.63
Figure 2: Feasibility and accuracy comparison of five selected LLMs on each individual task. The circles represent performance levels, progressing outward from 20% to 100% in increments of 20%.
In Table 1, we present the average performance of ten LLMs across four polynomial-time and six NP-complete tasks, assessed on both small and large graphs. As expected, GPT-4o consistently ranks as the top performer across most metrics and settings, with Llama3-70b closely following as the premier open-source model. Performance generally declines for all LLMs as graph sizes increase and tasks transition from polynomial-time to NP-complete, expanding the search space significantly. GPT-4o, despite its capabilities, struggles notably with NP-complete tasks on large graphs, achieving only a 5.93% accuracy rate. It indicates the most challenging scenario in GraphArena is still much larger than current LLMs’ upper bound.
GraphArena also reveals a significant performance gap between LLMs with different parameter numbers. For example, in polynomial-time tasks, the disparity between Llama3-70b and Llama3-8b is notable—61.2% vs. 28.6% on small graphs and 31.6% vs. 9.4% on large graphs. This gap is significantly wider than in other benchmarks, such as GSM8K (93.0% vs. 79.6%) and GPQA (39.5% vs. 34.2%). A similar trend is observed in the Qwen series. Moreover, models with fewer parameters are more prone to hallucination, with ratios of 78.2% for Llama3-8b, 91.2% for Qwen1.5-7b, and 84.0% for Gemma-7b, indicating that GraphArena’s challenges necessitate more advanced reasoning capabilities that larger LLMs are better equipped to manage.
To closely examine model performance on individual tasks, we analyzed the feasibility and accuracy of five selected LLMs—Claude3-haiku, Deepseek-V2, GPT-4o, Llama3-70b, and Mixtral-7x8b—as shown in Figure 2. Results for the remaining five LLMs are demonstrated in Figure 5 in the appendix. The general trends align with Table 1, showing decreased performance in larger graphs and more complex tasks. A detailed analysis shows that model performance varies substantially across different tasks. For instance, tasks such as Diameter, MCS, and MVC demonstrate notably low feasibility. This is attributed to their complex verification processes: verifying Diameter requires solving a shortest path problem, while verifying MCS involves graph isomorphism. Regarding accuracy, NP-complete tasks generally perform worse than polynomial-time tasks. Notably, the Diameter task shows the poorest performance among polynomial-time tasks due to its demand for a comprehensive global search. This indicates that task complexity significantly impacts accuracy.
Figure 3: The influence of graph size on hallucination probability for the Maximum Independent Set, Graph Diameter, and Connected Component tasks.
Figure 4: The influence of chain-of-thoughts with k-shot step-by-step demonstrations on the feasibility and accuracy of model performance for the Graph Diameter and Connected Component tasks.
3.2 In-depth Analysis of Hallucination
This section delves into the phenomenon of hallucination in LLMs, specifically aiming to address two critical questions: (1) What are the underlying causes of hallucination in GraphArena? (2) Are existing strategies able to mitigate the issue?
Our primary findings suggest that the hallucination ratio is linked with the number of model parameters, the complexity of its feasibility checks, and the scale of the graphs. To further illustrate this, we have plotted the hallucination probability against varying node sizes for three selected tasks in Figure 3. Additional data for seven other tasks are detailed in Figures 6 and 7 in the appendix. These plots reveal a significant, almost linear increase in hallucination probability as node size grows from 5 to 30. Notably, for GPT-4o on the Diameter task, the hallucination probability rises from 18% with a node size of 5 to 80% with a node size of 30, underscoring that larger graph sizes and their associated challenges in long text and exploding search space are major contributors to hallucination.
Chain-of-Thought (CoT) Prompting. Several strategies have been proposed to mitigate hallucinations in LLMs [9, 34, 87, 77]. We explore the effectiveness of a representative approach, CoT prompting, which encourages the model to generate intermediate reasoning steps towards a conclusion [78, 26, 52]. We design manually crafted examples for the Graph Diameter and Connected Component tasks, assessing the impact of using zero to four CoT examples. As shown in figure 4, a general enhancement in model accuracy is observed. Specifically, the performance of Deepseek-V2 notably improves after incorporating just one CoT example for the Component task. Nonetheless, Claude3-haiku and Mixtral-7x8b exhibit only marginal improvements. This suggests that, while CoT prompting contributes to reducing errors, it remains insufficient to fully resolve hallucinations.
Graph-aware LLMs. Several methods have been developed to fine-tune LLMs using graph-specific datasets, aiming to create graph-aware LLMs capable of understanding graph structures and solving basic graph tasks. To investigate whether this can help alleviate the issue of hallucinations, we test GraphArena on GraphWiz [17], an open-source model derived from Llama2-7b and fine-tuned using Direct Preference Optimization [61] across nine graph problems. While GraphWiz shows proficiency on the tasks it was initially trained on, it struggles on the GraphArena benchmark, with accuracies only ranging from 0.2% to 1.2% across four small-scale polynomial-time tasks. This can be attributed primarily to two factors: the absence of task overlap between GraphArena and the GraphWiz training corpus, and the model’s tendency to rely on pattern recognition rather than problem understanding.
Table 2: Performance comparison of LLMs and classic solvers on TSP for both small and large graphs. The table shows the percentage of problems where LLMs wins, ties, or loses against three classic solvers: Random, Greedy, and Christofides.
Travelling Salesman Problem (Small Graphs)
Solver Random Greedy Christofides
LLM win tie lose win tie lose win tie lose
GPT-4o 67.4 16.2 16.4 20.6 43.4 36.0 15.4 33.0 51.6
GPT-3.5 58.4 13.6 28.0 13.0 30.2 56.8 7.4 24.0 68.6
Claude3-haiku 54.8 14.4 30.8 15.4 25.2 59.4 9.4 21.0 69.6
Llama3-70b 59.8 18.4 21.8 12.0 36.6 51.4 8.4 23.8 67.8
Llama3-8b 43.0 13.6 43.4 12.2 15.4 72.4 6.4 15.6 78.0
Qwen1.5-72b 49.4 16.0 34.6 13.2 25.0 61.8 7.0 22.8 70.2
Qwen1.5-7b 33.6 11.2 55.2 10.8 14.2 75.0 6.0 14.8 79.2
Deepseek-V2 71.0 15.4 13.6 21.0 42.0 37.0 15.8 32.2 52.0
Mixtral-7x8b 43.6 16.8 39.6 10.6 22.6 66.8 6.0 18.0 76.0
Gemma-7b 27.6 15.0 57.4 8.8 13.8 77.4 5.0 13.0 82.0
Travelling Salesman Problem (Large Graphs)
GPT-4o 82.6 0.0 17.4 9.2 2.8 88.0 5.4 0.4 94.2
GPT-3.5 66.8 0.0 33.2 0.4 0.0 99.6 0.2 0.0 99.8
Claude3-haiku 72.6 0.0 27.4 1.0 0.0 99.0 0.6 0.0 99.4
Llama3-70b 79.6 0.0 20.4 1.4 0.0 98.6 0.6 0.0 99.4
Llama3-8b 29.8 0.0 70.2 0.2 0.0 99.8 0.0 0.0 100.0
Qwen1.5-72b 26.4 0.0 73.6 0.0 0.0 100.0 0.0 0.0 100.0
Qwen1.5-7b 23.0 0.0 77.0 0.0 0.0 100.0 0.0 0.0 100.0
Deepseek-V2 67.2 0.0 32.8 4.4 0.2 95.4 2.2 0.0 97.8
Mixtral-7x8b 23.6 0.0 76.4 0.0 0.0 100.0 0.0 0.0 100.0
Gemma-7b 3.8 0.0 96.2 0.0 0.0 100.0 0.0 0.0 100.0
3.3 Comparison with Classic Solvers
To better understand the performance of LLMs on NP-complete tasks, we compare them with several classic solvers on the Travelling Salesman Problem (TSP) task. We consider the following three solvers: (1) Random: randomly generates a feasible solution, (2) Greedy: selects a node at each step that is not already in the cycle and whose connection to the previous node adds the least cost to the cycle, and (3) Christofides: computes a 3/2-approximation of the TSP in a complete undirected graph using the Christofides algorithm [21], which is considered a better heuristic algorithm compared to Greedy. For each solver, the percentage of problems where the LLM wins, ties, or loses against the classic solvers is reported in Table 2.
In the context of small graphs, GPT-4o demonstrates a strong performance with a win rate of 67.4% against the Random solver, 20.6% against the Greedy solver, and 15.4% against the Christofides solver. Other models such as GPT-3.5 and Claude3-haiku show competitive performance but with slightly lower win rates. Smaller Models like Llama3-8b and Gemma-7b exhibit higher loss rates, indicating less effectiveness in solving TSP compared to the classic solvers.
For large graphs, GPT-4o has an even better win rate of 82.6% against the Random solver, but its performance drops significantly against the Greedy and Christofides solvers. Most other LLMs show similar trends. However, GPT-4o can still win in 27 problems compared to the Christofides solver, highlighting the potential of LLMs to provide strong heuristics in combinatorial optimization tasks.
4 Related Work
LLM Evaluation. Early evaluations of text generation focused on fundamental skills, including similarity measures such as ROUGE [46], METEOR [8], and BLEU [58] for machine translation [47, 64]. They also included inherent scores like Diversity [41] and Perplexity [33] for open-domain conversations [86, 43], and factual measures such as F-score and exact match for machine reading comprehension [62, 14] and question answering [90, 38]. Recently, numerous benchmarks have employed more advanced tests to assess the high-level capabilities of LLMs [30, 22, 50, 48, 15, 40, 74, 16, 71]. For instance, the Massive Multitask Language Understanding benchmark [29] covers 57 subjects across STEM, the humanities, the social sciences, and more. The Beyond the Imitation Game Benchmark [65] includes over 200 tasks, ranging from language understanding to interaction and behavior. Human preference remains a critical aspect of LLM evaluation [88, 42, 20]. Chatbot Arena [20] employs a pairwise comparison approach and leverages input from a diverse user base on crowd-sourcing. Our proposed GraphArena benchmark aims to evaluate a wide range of reasoning capabilities of LLMs on polynomial-time and NP-complete graph Computational problems, offering a valuable supplement to the existing benchmarks.
LLM on Graphs. The synergy between graphs and LLMs has sparked considerable interest due to their bi-directional benefits: integrating graph-based approaches can enhance the reasoning abilities of LLMs, enabling them to tackle complex logical tasks such as mathematical problem-solving [85], code generation [28], and knowledge reasoning [83, 57], among others. Conversely, LLMs offer powerful capabilities to augment graph analytics and learning, particularly for predictive tasks in text-attributed networks [25, 75, 18, 19, 70]. Compared with machine learning tasks on graphs such as node classification and anomaly detection [68, 69], graph computational problems pose more challenges for LLMs as they require a deeper understanding of structural information and long-range multi-step reasoning [51, 76, 67, 12]. GraphQA [25] and NLGraph [75] are two early benchmarks in this domain, focusing on basic graph problems using small-scale synthetic graphs. VisionGraph [45] and GITA [79] introduce multimodal graph reasoning tasks, which require LLMs to reason over both text and image modalities related to the graph. GraphWiz [17] introduces an instruction-tuning dataset designed to equip language models with the ability to tackle graph problems using explicit reasoning paths.
5 Conclusion
This paper introduces GraphArena, a benchmarking tool designed to evaluate the reasoning capabilities of LLMs through graph computational challenges. GraphArena features a realistic graph collection, carefully selected tasks, and a rigorous evaluation framework. Our evaluation of ten LLMs across 10,000 GraphArena problems reveals persistent hallucination issues, especially with larger graphs, more complex tasks, and models with fewer parameters. Despite implementing advanced strategies such as chain-of-thought prompting and domain-specific fine-tuning, these issues remain unresolved. Future efforts include investigating LLMs’ ability to write code, integrating multi-modality to better address these challenges, and exploring new techniques to mitigate hallucinations and improve reasoning capabilities.
Appendix A Additional Experiments
We provide detailed statistics and examples to further explain the tasks and performance evaluations discussed in the main text. Table 3 presents comprehensive statistics for each task in GraphArena, including the average and maximum number of nodes edges, and text length counted by number of characters. Figure 5 offers a feasibility and accuracy comparison of the remaining five LLMs on each individual task, complementing the analysis in Figure 2. Figures 6 and 7 illustrate the influence of graph size on hallucination probability for the remaining seven tasks, as the complement of Figure 3. Tables LABEL:table:COT_Component_prompt through LABEL:table:TSP_prompt provide detailed examples of prompts for each task. For the Connected Component and Graph Diameter tasks, chain-of-thought (CoT) prompting with step-by-step solution demonstrations is provided.
Table 3: Statistics of problems in each task of GraphArena, including the average and maximum number of nodes ($V$), edges ($E$), and text length ($T$).
Problems avg. $\left|V\right|$ max. $\left|V\right|$ avg. $\left|E\right|$ max. $\left|E\right|$ avg. $\left|T\right|$ max. $\left|T\right|$
Common Neighbor 19.4 50 68.6 409 4323.5 18014
Shortest Distance 19.6 50 36.8 209 5639.5 23731
Connected Component 13.5 30 14.4 122 1785.4 5503
Graph Diameter 13.4 30 22.0 128 3741.3 11783
Maximum Clique Problem 13.2 30 39.1 154 3025.0 7921
Maximum Independent Set 12.8 30 19.5 141 1902.8 6182
Minimum Vertex Cover 14.1 30 23.3 151 2178.4 6429
Maximum Independent Set 9.2 20 8.9 22 1025.1 1187
Graph Edit Distance 9.4 20 9.2 23 1643.7 2142
Travelling Salesman Problem 9.5 20 50.6 190 1806.1 4741
Figure 5: Feasibility and accuracy comparison of the remaining five LLMs in Figure 2 on each individual task.
Figure 6: The influence of graph size on hallucination probability for the tasks of Distance, Neighbor, MCP, and MVC.
Figure 7: The influence of graph size on hallucination probability for the tasks of GED, TSP, and MCS.
Table 4: An example of CoT prompting with step-by-step demonstrations on the Component task.
CoT Prompts of the Connected Component Task. (1-shot)
You are required to identify all connected components in the given social network and output one representative node from each component.
Within a connected component, any node can be reached from any other node through the edges in the graph. Different connected components are isolated from each other.
**Example**
- Names in the network: Michelle Nelson, Kevin Brooks, Eric Davis, Toni Suarez, Teresa Hughes
- Friendship connections: Michelle Nelson to Teresa Hughes, Kevin Brooks to Eric Davis, Eric Davis to Toni Suarez
The answer includes one representative element from each connected component in the given social network: In this social network, we have the following names and connections: Michelle Nelson, Kevin Brooks, Eric Davis, Toni Suarez, and Teresa Hughes. The connections are: Michelle Nelson to Teresa Hughes, Kevin Brooks to Eric Davis, and Eric Davis to Toni Suarez. From these connections, we can identify two isolated connected components. The first component includes Michelle Nelson and Teresa Hughes, and the second component includes Kevin Brooks, Eric Davis, and Toni Suarez. By selecting one representative node from each connected component, we have Teresa Hughes from the first component and Kevin Brooks from the second component. Therefore, the answer is [Teresa Hughes, Kevin Brooks].
**Problem to Solve**
- Names in the network: Michelle Nelson, Kevin Brooks, Eric Davis, Toni Suarez, Teresa Hughes
- Friendship connections: Michelle Nelson to Teresa Hughes, Kevin Brooks to Eric Davis, Eric Davis to Toni Suarez
Identify all connected components in this network. Note that for each connected component, you should only output one of its nodes. Present your answer in the following format: [UserA, UserB, UserC, UserD, …]
Table 5: An example of CoT prompting with step-by-step demonstrations on the Diameter task.
CoT Prompts of the Graph Diameter Task. (1-shot)
You are required to calculate the diameter of an undirected knowledge graph.
The diameter of a graph is the maximum distance between any pair of nodes in the graph. To compute this, you need to find the shortest path between all pairs of nodes and then determine the maximum length of these shortest paths.
**Example**
-Entities in this knowledge graph: SeineetMarne, MontereauFaultYonne, Vittorio De Sica, SaintFiacre SeineetMarne, France
- The relationships between these entities are as follows: SeineetMarne is connected to SaintFiacre SeineetMarne via the relationship department. SeineetMarne is connected to MontereauFaultYonne via the relationship department. SeineetMarne is connected to France via the relationship country. MontereauFaultYonne is connected to France via the relationship country. Vittorio De Sica is connected to France via the relationship deathPlace. SaintFiacre SeineetMarne is connected to France via the relationship country.
To calculate the diameter of this undirected knowledge graph, we first need to understand the connections and the structure of the graph. The entities and their relationships are as follows: SeineetMarne is connected to SaintFiacre SeineetMarne via the relationship ’department’, SeineetMarne is connected to MontereauFaultYonne via the relationship ’department’, SeineetMarne is connected to France via the relationship ’country’, MontereauFaultYonne is connected to France via the relationship ’country’, Vittorio De Sica is connected to France via the relationship ’deathPlace’, and SaintFiacre SeineetMarne is connected to France via the relationship ’country’. The graph structure can be visualized as follows: SeineetMarne is connected to SaintFiacre SeineetMarne, MontereauFaultYonne, and France. MontereauFaultYonne and SaintFiacre SeineetMarne are also connected to France, and France is connected to Vittorio De Sica. By analyzing the shortest paths between all pairs of nodes, we find that SeineetMarne to SaintFiacre SeineetMarne, SeineetMarne to MontereauFaultYonne, SeineetMarne to France, MontereauFaultYonne to France, SaintFiacre SeineetMarne to France, and France to Vittorio De Sica all have a direct connection with 1 edge. The longest shortest paths are from SeineetMarne, MontereauFaultYonne, and SaintFiacre SeineetMarne to Vittorio De Sica via France, each with 2 edges. Therefore, the diameter of this knowledge graph is 2, and one of the paths that represent this diameter is [SeineetMarne, France, Vittorio De Sica].
**Problem to Solve**
- Entities in this knowledge graph: PlayStation 2, Buzz! Junior: Jungle Party, Tekken (video game), Buzz!, PlayStation Network
- The relationships between these entities are as follows:
PlayStation 2 is connected to Tekken (video game) via the relationship computingPlatform. PlayStation 2 is connected to Buzz! Junior: Jungle Party via the relationship computingPlatform. Buzz! Junior: Jungle Party is connected to Buzz! via the relationship series. Buzz! Junior: Jungle Party is connected to PlayStation Network via the relationship computingPlatform. Tekken (video game) is connected to PlayStation Network via the relationship computingPlatform.
Please determine the diameter of this network and output the corresponding path in the following format: [Entity1, Entity2, …, EntityN].
Table 6: An example of the Common Neighbor task.
Prompts of the Common Neighbor Task. (1-shot)
Your task is to find the common neighbors of two nodes in an undirected academic network. In this network, nodes represent authors and edges represent research collaborations.
**Example**
- Authors in the network: Marie-Francine Moens, Stefanie Brüninghaus, Henry Prakken, David W. Aha, Kevin D. Ashley, Ronald Prescott Loui
- Research collaborations between these authors: Marie-Francine Moens and Kevin D. Ashley, Marie-Francine Moens and Stefanie Brüninghaus, Stefanie Brüninghaus and David W. Aha, Stefanie Brüninghaus and Henry Prakken, Stefanie Brüninghaus and Kevin D. Ashley, Stefanie Brüninghaus and Ronald Prescott Loui, Henry Prakken and Kevin D. Ashley, Henry Prakken and Ronald Prescott Loui, David W. Aha and Kevin D. Ashley, Kevin D. Ashley and Ronald Prescott Loui.
Common neighbors between Marie-Francine Moens and Stefanie Brüninghaus: [Kevin D. Ashley] **Problem to Solve**
- Authors in the network: Sang Wan Lee, Hyoyoung Jang, Z. Zenn Bien, Zeungnam Bien
- Research collaborations between these authors: Sang Wan Lee and Zeungnam Bien, Sang Wan Lee and Z. Zenn Bien, Sang Wan Lee and Hyoyoung Jang, Hyoyoung Jang and Zeungnam Bien, Z. Zenn Bien and Zeungnam Bien.
Please identify the common neighbors of Sang Wan Lee and Hyoyoung Jang in this network. Present your answer in the following format: [AuthorA, AuthorB, AuthorC, AuthorD, …].
Table 7: An example of the Shortest Distance task.
Prompts of the Shortest Distance Task. (1-shot)
Your task is to identify the shortest path between two specified entities in an undirected knowledge graph, minimizing the number of hops.
**Example**
- Entities in this knowledge graph: Chordata, Animalia, Sloggett’s vlei rat, Mammalia
- The relationships between these entities are as follows:
Chordata is connected to Sloggett’s vlei rat via the relationship phylum. Animalia is connected to Sloggett’s vlei rat via the relationship kingdom. Sloggett’s vlei rat is connected to Mammalia via the relationship class.
One shortest path between Animalia and Sloggett’s vlei rat is: [Animalia, Sloggett’s vlei rat]
**Problem to Solve**
- Entities in this knowledge graph: United States, Post-hardcore, Head Wound City, Texas, Cody Votolato
- The relationships between these entities are as follows:
United States is connected to Texas via the relationship country. United States is connected to Cody Votolato via the relationship birthPlace. Post-hardcore is connected to Cody Votolato via the relationship genre. Head Wound City is connected to Cody Votolato via the relationship associatedMusicalArtist. Texas is connected to Cody Votolato via the relationship birthplace.
Please determine the shortest path between Texas and United States in this network.Submit your answer in the format: [Entity1, Entity2, …, EntityN], where Entity1 and EntityN are the specified start and end entities, and Entity2 through EntityN-1 are the intermediate entities on the shortest path.
Table 8: An example of the Graph Edit Distance task.
Prompts of the GED Task. (1-shot)
You are required to solve the Graph Edit Distance problem between two molecules. Each edit operation (adding or deleting an edge, adding or deleting an isolated node, or relabeling a node) has an identity cost. Your objective is to establish a mapping between the atom IDs from Molecule A to Molecule B, ensuring that each atom ID in Molecule A corresponds to exactly one atom ID in Molecule B. The mapping corresponds to the minimum edit cost between the two graphs.
**Example**
Molecule A:
- Atoms: N (atom 0), O (atom 1), Si (atom 2), O (atom 3), O (atom 4).
- Bonds: 0-1, 1-2, 2-3, 2-4.
Molecule B:
- Atoms: F (atom 0), B (atom 1), C (atom 2), N (atom 3), Br (atom 4).
- Bonds: 0-1, 1-2, 1-4, 2-3.
One optimal node mapping: [3, 2, 1, 0, 4].
**Problem to Solve**
You are given the following two molecules:
Molecule A:
Atoms: N (atom 0), C (atom 1), N (atom 2), F (atom 3).
Bonds: 0-1, 1-2, 1-3.
Molecule B:
Atoms: O (atom 0), C (atom 1), F (atom 2), F (atom 3).
Bonds: 0-1, 1-2, 1-3.
Represent the node mapping as a list of integers, where the position in the list corresponds to the atom ID in Molecule A and the value at that position indicates the corresponding atom ID in Molecule B.
For instance, if atom 0 in Molecule A corresponds to atom 1 in Molecule B, atom 1 in Molecule A corresponds to atom 0 in Molecule B, and atom 2 remains unchanged, the mapping would be represented as [1, 0, 2, …].
Table 9: An example of the Maximum Clique Problem.
Prompts of the MCP Task. (1-shot)
You are required to solve the Maximum Clique Problem for an undirected academic network. In this network, nodes represent authors and edges represent research collaborations. Your objective is to find the largest subset of nodes such that every pair of vertices in this subset is connected by an edge.
**Example**
- Authors in the network: Keng Peng Tee, Veit Hagenmeyer, Bartosz Käpernick, Karl Henrik Johansson, Darryl DeHaan, James B. Rawlings, Andreas Kugi, Knut Graichen, Tilman Utz, Christian Ebenbauer.
- Research collaborations between these authors: Keng Peng Tee and James B. Rawlings, Keng Peng Tee and Darryl DeHaan, Veit Hagenmeyer and Knut Graichen, Bartosz Käpernick and Andreas Kugi, Bartosz Käpernick and Knut Graichen, Karl Henrik Johansson and Christian Ebenbauer, Darryl DeHaan and Knut Graichen, Darryl DeHaan and Christian Ebenbauer, James B. Rawlings and Knut Graichen, James B. Rawlings and Christian Ebenbauer, Andreas Kugi and Knut Graichen, Knut Graichen and Tilman Utz.
One Maximum Clique: [Knut Graichen, Bartosz Käpernick, Andreas Kugi].
**Problem to Solve**
- Authors in the network: Manfred Schmidt-Schauss, David Sabel, Manfred Schmidt-Schauß, Guillem Godoy.
- Research collaborations between these authors: Manfred Schmidt-Schauss and David Sabel, Manfred Schmidt-Schauss and Manfred Schmidt-Schauß, Manfred Schmidt-Schauss and Guillem Godoy, David Sabel and Manfred Schmidt-Schauß, Manfred Schmidt-Schauß and Guillem Godoy.
Identify the clique with the maximum number of authors in this network. Present your answer in the following format: [AuthorA, AuthorB, AuthorC, AuthorD, …].
Table 10: An example of the Maximum Common Subgraph Task.
Prompts of the MCS Task. (1-shot)
You are required to solve the Maximum Common Subgraph problem. Your goal is to identify the common subgraph with the maximum number of atoms shared between the two molecules.
**Example**
Molecule A consists of 8 atoms with the following 9 bonds: 0-1, 0-6, 1-2, 2-3, 3-4, 3-7, 4-5, 5-6, 5-7.
Molecule B consists of 7 atoms with the following 7 bonds: 0-1, 1-2, 1-4, 2-3, 3-4, 3-6, 4-5.
One max common subgraph: [2, 3, 4, 5, 7, 6], [0, 1, 2, 3, 4, 6].
**Problem to Solve**
You are given the following two molecules:
Molecule A consists of 4 atoms with the following 3 bonds: 0-1, 1-2, 2-3. Molecule B consists of 4 atoms with the following 3 bonds: 0-1, 1-2, 1-3.
Provide the indices of the atoms in the common subgraph for each molecule in the following format: [Node indices in molecular A], [Node indices in molecular B].
For example, if the common subgraph is the subgraph of atom 1, 2, 3 in molecule A and the subgrah of atom 2, 3, 4 in molecule B, you should answer: [1, 2, 3], [2, 3, 4].
Table 11: An example of the Maximum Independent Set task.
Prompts of the MIS Task. (1-shot)
Your task is to solve the Maximum Independent Set problem in the given social network. In this network, each node represents a user, and each edge represents a friendship connection. You need to identify the largest subset of users such that no two users in this subset are friends connected by an edge.
**Example**
- Users in the network: Melinda Vaughan, Mary Thornton, Jeremiah Griffith, Lisa Anderson, Alfred Powell.
- Fiendship connections: Melinda Vaughan and Jeremiah Griffith, Mary Thornton and Jeremiah Griffith, Jeremiah Griffith and Lisa Anderson, Jeremiah Griffith and Alfred Powell.
One Maximum Independent Set: [Melinda Vaughan, Lisa Anderson, Alfred Powell, Mary Thornton].
**Problem to Solve**
- Users in the network: William Lawson, Daniel Shelton, Michelle Lewis, Julie Hayes.
- Friendship connections: William Lawson and Daniel Shelton, William Lawson and Julie Hayes, Michelle Lewis and Julie Hayes.
Identify the Maximum Independent Set of this network and present your answer in the following format: [UserA, UserB, UserC, UserD, …].
Table 12: An example of the Minimum Vertex Cover task.
Prompts of the MVC Task. (1-shot)
Your task is to solve the Minimum Vertex Cover problem in the given social network. In this network, each node represents a user, and each edge represents a friendship connection. You need to identify the smallest subset of users such that every friendship connection has at least one user from this subset.
**Example**
- Users in the network: Julie Harris, David Torres, Vanessa Parker, Shawn Barnett, Karl Dean.
- Fiendship connections: Julie Harris and Vanessa Parker, Julie Harris and David Torres, Julie Harris and Shawn Barnett, Julie Harris and Karl Dean, David Torres and Vanessa Parker, David Torres and Shawn Barnett, David Torres and Karl Dean, Vanessa Parker and Shawn Barnett, Shawn Barnett and Karl Dean.
One Minimum Vertex Cover: [Julie Harris, David Torres, Shawn Barnett].
**Problem to Solve**
- Users in the network: Pamela Haynes, Kyle Meadows, Adam Nichols, Anna Lowery, Heather Dixon, Matthew Lee, Elizabeth Wood, Stephen Hess.
- Friendship connections: Pamela Haynes and Stephen Hess, Kyle Meadows and Matthew Lee, Kyle Meadows and Stephen Hess, Kyle Meadows and Adam Nichols, Adam Nichols and Stephen Hess, Adam Nichols and Heather Dixon, Anna Lowery and Stephen Hess, Heather Dixon and Stephen Hess, Matthew Lee and Stephen Hess, Elizabeth Wood and Stephen Hess.
Identify the Minimum Vertex Cover of this network and present your answer in the following format: [UserA, UserB, UserC, UserD, …].
Table 13: An example of the Travelling Salesman Problem.
Prompts of the TSP Task. (1-shot)
You are required to solve the Travelling Salesman Problem for an undirected flight route network. Your objective is to determine the shortest possible route that visits each of the listed airports exactly once and returns to the starting point.
**Example**
- Airports to visit: VPY, AAE, BGA, YWB.
- Travel distances (in kilometers) between each pair of airports:
VPY to YWB: 16285; VPY to AAE: 9488; VPY to BGA: 13255; AAE to YWB: 7807; AAE to BGA: 9332; BGA to YWB: 6575.
One shortest route: [VPY, AAE, YWB, BGA, VPY].
**Problem to Solve**
- Airports to visit: YNT, TEB, CFC, VIE, NSH, ROV.
- Travel distances (in kilometers) between each pair of airports:
YNT to VIE: 7962; YNT to ROV: 6644; YNT to NSH: 6176; YNT to CFC: 18668; YNT to TEB: 12263; TEB to VIE: 6987; TEB to ROV: 8529; TEB to NSH: 10389; TEB to CFC: 8846; CFC to VIE: 10716; CFC to ROV: 12244; CFC to NSH: 13193; VIE to ROV: 1736; VIE to NSH: 3406; NSH to ROV: 3000.
Please calculate the shortest tour and format your answer as follows: [Airport A, Airport B, Airport C, …, Airport A]
Identify the Minimum Vertex Cover of this network and present your answer in the following format: [UserA, UserB, UserC, UserD, …].